# SynGraph: A Dynamic Graph-LLM Synthesis Framework for Sparse Streaming User Sentiment Modeling

### v1

@@ Abstract
Xin Zhang1, Qiyu Wei1, Yingjie Zhu2, Linhai Zhang3, Deyu Zhou4, Sophia Ananiadou1 1The University of Manchester  2Harbin Institute of Technology 3King’s College London  4Southeast University {xin.zhang-41@postgrad., qiyu.wei@postgrad., sophia.ananiadou@}manchester.ac.uk zhuyj@stu.hit.edu.cn, linhai.zhang@kcl.ac.uk, d.zhou@seu.edu.cn
@@ 1 Introduction
User reviews on e-commerce platforms provide valuable insights into customer preferences, product performance, and user satisfaction. These reviews reflect users’ momentary sentiments and exhibit temporal dynamics as users interact with products over time. Traditional sentiment analysis methods, however, predominantly operate in static contexts, where individual reviews are treated as independent data points Tang et al. (2015); Chen et al. (2016); Kim and Srivastava (2007); Mittal et al. (2022). This paradigm overlooks the inherent temporal dependencies and evolving nature of user sentiments, which are crucial for capturing the full spectrum of user behavior in real-world scenarios. For instance, as shown in Figure 1, a user’s sentiment rating may initially be highly positive (e.g., “I really enjoyed this book”) but gradually shift to neutral (e.g., “I was surprised that I liked this book as much as I did”) in subsequent interactions. These transitions highlight the dynamic relationship between sentiment changes and textual cues, require models to learn sentiment change patterns from historical reviews, and leverage this understanding to predict future sentiment ratings.
@@
As a solution to these limitations, the task of “Sentiment Analysis on Streaming Reviews” has been proposed, aiming to model the temporal evolution of user sentiments across sequential time windows and predict user future sentiments Zhang et al. (2023); Wu et al. (2023). This task emphasizes the interplay between numerical sentiment scores and their associated textual content over time. However, this task also introduces significant challenges, primarily due to the problem of data sparsity Guo (2013); Du et al. (2022). Previous studies have commonly introduced graph structure information to supplement sparse user data in static scenarios (Zhou et al., 2023; Wang et al., 2023; Chen et al., 2022). Despite this, in streaming scenarios, the inclusion of graph structure information further complicates the classification and discussion of sparse users Xu et al. (2024). As shown in Figure 1, some users leave reviews at irregular and widely spaced intervals, making it challenging for temporal models to capture sequential dependencies (User A: temporal sparsity). Others have limited social connections, resulting in a lack of neighbor information to support sentiment analysis (User B: spatial sparsity). Additionally, some users exhibit both temporal and spatial sparsity, further complicating predictive modeling (User C: combined sparsity). Existing methods fail to effectively address these sparsity issues, struggling to model temporal dependencies and spatial interactions.

### v2

@@ Abstract
Xin Zhang1, Qiyu Wei1, Yingjie Zhu2, Linhai Zhang3, Deyu Zhou4, Sophia Ananiadou1 1The University of Manchester  2Harbin Institute of Technology 3King’s College London  4Southeast University {xin.zhang-41, qiyu.wei}@postgrad.manchester.ac.uk sophia.ananiadou@manchester.ac.uk, zhuyj@stu.hit.edu.cn linhai.zhang@kcl.ac.uk, d.zhou@seu.edu.cn
@@ 1 Introduction
User reviews on e-commerce platforms provide valuable insights into customer preferences, product performance, and user satisfaction. These reviews reflect users’ momentary sentiments and exhibit temporal dynamics as users interact with products over time. Traditional sentiment analysis methods, however, predominantly operate in static contexts, where individual reviews are treated as independent data points Tang et al. (2015); Chen et al. (2016); Kim and Srivastava (2007); Mittal et al. (2022); DBLP:journals/kbs/ZhangZZ25. This paradigm overlooks the inherent temporal dependencies and evolving nature of user sentiments, which are crucial for capturing the full spectrum of user behavior in real-world scenarios. For instance, as shown in Figure 1, a user’s sentiment rating may initially be highly positive (e.g., “I really enjoyed this book”) but gradually shift to neutral (e.g., “I was surprised that I liked this book as much as I did”) in subsequent interactions. These transitions highlight the dynamic relationship between sentiment changes and textual cues, require models to learn sentiment change patterns from historical reviews, and leverage this understanding to predict future sentiment ratings.
@@
As a solution to these limitations, the task of “Sentiment Analysis on Streaming Reviews” has been proposed, aiming to model the temporal evolution of user sentiments across sequential time windows and predict user future sentiments Zhang et al. (2023); Wu et al. (2023). This task emphasizes the interplay between numerical sentiment scores and their associated textual content over time. However, this task also introduces significant challenges, primarily due to the problem of data sparsity Guo (2013); Du et al. (2022); DBLP:journals/corr/abs-2403-06139. Previous studies have commonly introduced graph structure information to supplement sparse user data in static scenarios (Zhou et al., 2023; Wang et al., 2023; Chen et al., 2022). Despite this, in streaming scenarios, the inclusion of graph structure information further complicates the classification and discussion of sparse users Xu et al. (2024). As shown in Figure 1, some users leave reviews at irregular and widely spaced intervals, making it challenging for temporal models to capture sequential dependencies (User A: temporal sparsity). Others have limited social connections, resulting in a lack of neighbor information to support sentiment analysis (User B: spatial sparsity). Additionally, some users exhibit both temporal and spatial sparsity, further complicating predictive modeling (User C: combined sparsity). Existing methods fail to effectively address these sparsity issues, struggling to model temporal dependencies and spatial interactions.
@@ 6 Conclusion
Acknowledgement
We would like to thank the anonymous reviewers for their insightful comments and suggestions. This work was supported by the British Heart Foundation Manchester Research Excellence Award (RE/24/130017). We also acknowledge the CSF3 at the University of Manchester for providing GPU resources. Xin is supported by the UoM-CSC Joint Scholarship.
@@ Appendix B Prompt Templates
Appendix C Synthetic-Data Proportion Experiments
We conduct a series of controlled experiments on three datasets (Appliances, Gift_Cards, Magazine_Subscriptions) using the DC-DGNN framework, in which synthetic data is inserted at six stages: 0% (NoSynth), Front 20%, Front 40%, Front 60%, Front 80%, and 100% (FullSynth).
Appliances (approximately 100 users)
In this extremely sparse setting, adding a small amount of synthetic data (Front 20%) yields negligible changes in Accuracy or F1 and may introduce minor fluctuations due to noise. As the synthetic data ratio increases most metrics improve but Accuracy at Front 80% slightly dips, which indicates that moderate augmentation often outperforms both no augmentation and full augmentation in very small datasets (see Table 5).
Gift_Cards and Magazine_Subscriptions (larger scale)
Both datasets show a generally monotonic increase in Accuracy and F1 as the synthetic data ratio increases, which suggests that when ample real data is available the negative impact of noise is reduced and the model benefits consistently from additional synthetic samples (see Tables 6 and 7).
Based on these results, we conclude that for very small and sparse datasets such as Appliances with only approximately 100 users moderate synthetic data ratios (for example 40 to 60%) achieve the best balance between performance improvement and noise. Therefore, the synthetic data ratio and insertion stage must be calibrated on a per dataset basis.
@@ Appendix C Synthetic-Data Proportion Experiments
Variation Accuracy Precision Recall F1 MSE RMSE MAE
Appliances_0%_NoSynth 0.7143 0.2381 0.3333 0.2778 1.0476 1.0235 0.4762
Appliances_Front200.7143 0.2381 0.3333 0.2778 1.0476 1.0235 0.4762
Appliances_Front400.7619 0.4250 0.4778 0.4492 0.5238 0.7237 0.3333
Appliances_Front600.7619 0.4556 0.5389 0.4889 0.7619 0.8729 0.3810
Appliances_Front800.7143 0.4000 0.4556 0.4222 0.5714 0.7559 0.3810
Appliances_FullSynth 0.8571 0.5441 0.5833 0.5625 0.6667 0.8165 0.2857
Table 5: Results on the Appliances dataset under different synthetic data proportions.
Variation Accuracy Precision Recall F1 MSE RMSE MAE
GiftCards_0%_NoSynth 0.8754 0.2189 0.2500 0.2334 0.3064 0.5535 0.1717
GiftCards_Front200.8721 0.1780 0.1992 0.1880 0.1751 0.4184 0.1414
GiftCards_Front400.8721 0.4732 0.2897 0.3064 0.1481 0.3849 0.1347
GiftCards_Front600.8923 0.3787 0.3667 0.3706 0.1279 0.3577 0.1145
GiftCards_Front800.8990 0.7241 0.5086 0.5030 0.1212 0.3482 0.1077
GiftCards_FullSynth 0.8956 0.4384 0.5000 0.4671 0.1145 0.3383 0.1077
Table 6: Results on the Gift_Cards dataset under different synthetic data proportions.
Variation Accuracy Precision Recall F1 MSE RMSE MAE
Magazine_0%_NoSynth 0.7554 0.4290 0.4107 0.4016 0.7768 0.8814 0.3820
Magazine_Front200.7639 0.5851 0.4753 0.4830 0.5193 0.7206 0.3133
Magazine_Front400.7768 0.4126 0.4794 0.4370 0.5064 0.7116 0.3004
Magazine_Front600.8155 0.7670 0.6302 0.6210 0.3133 0.5597 0.2189
Magazine_Front800.8155 0.5576 0.5495 0.4991 0.3906 0.6249 0.2361
Magazine_FullSynth 0.7983 0.6879 0.5853 0.5385 0.4206 0.6485 0.2575
Table 7: Results on the Magazine_Subscriptions dataset under different synthetic data proportions.